Question 16 Not yet answered Points out of4.00 P Flag question If the probability that an event will happen is 0.27, what is the probability that the event wont happen? Select one: O a.0.83 O b.0.23 О С. 063 o d. 0.73 O e. 1.73 Time left 0:12:44 Type here to search PrtSen rs
Answer #1

Here the probability that an event will happen is 0.27.

Now the probability that event will not happen is

P(event will not happen)=1-P(event will happen)=1-0.27=0.73

ANS: d.0.73